Your instructor (Katie) believes that the average time of taking this test is equal to 140 minutes. A sample of 4 students was taken and the following test times were obtained. Assume the distribution of the population is normally distributed.

Sample: 150, 150, 180, 170 (Sample Std. Dev = 15)

(1) The critical value for this test at a 0.1 level of significance is:

a.

1.285, -1.285

b.

1.645, -1.645

c.

1.638, -1.638

d.

2.353, -2.353

(2) At a 0.1 level of significance, it can be concluded that the mean of the population is:

a.

significantly less than 140

b.

not significantly different than 140

c.

significantly greater than 140

d.

significantly different than 140
Answer #1

given data are:-

sample size (n) = 4

sample sd (s) = 15

sample mean (T) = (150+150+180+170)/4 = 162.5

a). The critical value for this test at a 0.1 level of significance is: -2.353 , 2.353

[ df = (n-1) = (4-1) = 3

t critical value for alpha=0.10, df = 3, both tailed test be = \pm 2.353 ]

b). significantly different than 140

[ explanation:-

hypothesis:-

H_{0}:\mu=140

H_{1}:\mu\neq140

the test statistic be:-

t=\frac{\bar{x}-\mu}{s/\sqrt{n}}=\frac{162.5-140}{15/\sqrt{4}}=\mathbf{3}

decision:-

t=3 > 2.353

we reject the null hypothesis . ]

*** if you have any doubts about the problem please mention it in the comment box... if satisfied please LIKE.